### v1

# Wave dynamics governing vortex breakdown in smooth Euler flows

Abstract.
We consider the three-dimensional incompressible Euler equations in a setting where a vortex is transported by a prescribed Lagrangian flow. We show that vortex breakdown is governed by wave dynamics generated by the underlying transport flow. The key idea is to avoid any singular integral representation of the pressure term and instead construct an effective Lagrangian coordinate system in which the associated Lie bracket vanishes identically.
Key words and phrases:
Euler equations, vortex breakdown, Frenet-Serret formulas
2010 Mathematics Subject Classification:
Primary 35Q31; Secondary 76B47; Tertiary 76B55
1. Introduction
One of the most important features of Navier–Stokes turbulence is that it is not random but is instead organized around vortex stretching dynamics. More precisely, recent direct numerical simulations of Navier–Stokes turbulence at sufficiently high Reynolds numbers [2, 3, 10, 11] have reported the existence of a hierarchical structure of vortex-stretching motions. In particular, Goto, Saito, and Kawahara [3] clearly observed that turbulence at sufficiently high Reynolds numbers in a periodic cube exhibits a self-similar hierarchy of antiparallel vortex-tube pairs, which is sustained by the generation of smaller-scale vortices through stretching within larger-scale strain fields. This observation was further investigated by Y–Goto–Tsuruhashi [14] and Tsuruhashi–Goto–Oka–Y [13] (see also [5, 6, 7] for related mathematical results). Following this line of research, attention naturally turns to the dynamics of vortex stretching and the mechanisms underlying the subsequent breakdown of vortices.
With this perspective in mind, in this paper we provide a mathematical analysis of vortex breakdown in a scenario where a vortex is transported by a prescribed Lagrangian flow. Let $\Phi$ denote the Lagrangian flow map associated with the incompressible Euler equations, written in Lagrangian form (omitting the explicit volume-preserving constraint) as follows:
(1) $\left(\partial_{t}^{2}\Phi(t,\cdot)\right)\circ\Phi^{-1}(t,x)=-\nabla p(t,x),\qquad t>0,\ x\in\mathbb{R}^{3}.$
Let $\ell=\bigcup_{s}\ell(s)\subset\mathbb{R}^{3}$ be a smooth curve representing the vortex axis. We now describe the time evolution of swirling particles around $\ell$. Let $z$ denote the arc-length parameter along the transported curve $\bigcup_{t}\Phi(t,\ell(s))$. More precisely, we define a reparametrization $t=t(z)$ (implicitly depending on $s$) by
(2) $\partial_{z}t=v(t)^{-1}\qquad v(t):=\left|\partial_{t}\Phi(t,\ell(s))\right|,$
so that
$\left|\partial_{z}\Phi(t(z),\ell(s))\right|=1.$
Note that this implies $\partial_{z}=v(t)^{-1}\partial_{t}$. Using the transported curve $\bigcup_{t}\Phi(t,\ell(s))$, we introduce a Lagrangian moving frame. First, we define the unit tangent vector
$\tau(z):=\partial_{z}\Phi(t(z),\ell(s)).$
The curvature $\kappa(z)$ and unit normal vector $n(z)$ are defined by
$\partial_{z}\tau(z)=\kappa(z)\,n(z).$
The torsion $T(z)$ and binormal vector $b(z)$ are then defined through the Frenet–Serret relation:
$\partial_{z}n(z)=-\kappa(z)\,\tau(z)+T(z)\,b(z).$
For any initial particle
$x=\ell(s)+Z(0,r_{1},r_{2})\,\tau(0)+r_{1}\,n(0)+r_{2}\,b(0),$
the flow map $\Phi(t,x)$ can be uniquely expressed (for sufficiently small $r_{1},r_{2}$) as
(3) $\begin{split}\Phi(t,x)&=:\Phi(t,s,r_{1},r_{2})\\ &=\Phi(t,\ell(s))+Z(t,r_{1},r_{2})\,\tau(t)+R_{1}(t,r_{1},r_{2})\,n(t)+R_{2}(t,r_{1},r_{2})\,b(t).\end{split}$
Here
$Z(0,r_{1},r_{2})=\alpha_{1}(0)r_{1}+\alpha_{2}(0)r_{2},\qquad Z(t,0,0)=0,$
and
$R_{1}(0,r_{1},r_{2})=r_{1},\qquad R_{2}(0,r_{1},r_{2})=r_{2},$
with $\alpha_{1}(0),\alpha_{2}(0)\in\mathbb{R}$. Since the Jacobian
$\frac{\partial(R_{1},R_{2})}{\partial(r_{1},r_{2})}$
is nonzero at $(r_{1},r_{2})=(0,0)$, the inverse function theorem implies that, for each fixed $t>0$ and for sufficiently small $|R_{1}|,|R_{2}|$, we may rewrite (3) in terms of $(R_{1},R_{2})$ as follows. Let $r_{1}=r_{1}(t,R_{1},R_{2})$ and $r_{2}=r_{2}(t,R_{1},R_{2})$ denote the corresponding inverse functions. Then, for any initial particle
$x=\ell(s)+r_{1}(t,R_{1},R_{2})\,n(0)+r_{2}(t,R_{1},R_{2})\,b(0),$
we have
(4) $\begin{split}\Phi(t,x)&=:\Phi(t,s,R_{1},R_{2})\\ &=\Phi(t,\ell(s))+Z\bigl(t,r_{1}(t,R_{1},R_{2}),r_{2}(t,R_{1},R_{2})\bigr)\,\tau(t)\\ &\quad+R_{1}\,n(t)+R_{2}\,b(t).\end{split}$
Note that $\Phi(t,s,R_{1},R_{2})\big|_{R_{1}=R_{2}=0}$ represents the Lagrangian transport of the vortex axis $\ell$. We now simplify the function $Z$ by the expansion
(5) $Z\bigl(t,r_{1}(t,R_{1},R_{2}),r_{2}(t,R_{1},R_{2})\bigr)=\alpha_{1}(t)R_{1}+\alpha_{2}(t)R_{2}+\mathcal{O}(R_{1}^{2}+R_{2}^{2}),$
and we interpret $Z$ as describing the time evolution of swirling particles around the vortex axis. The initial data $\alpha_{1}(0),\alpha_{2}(0)\in\mathbb{R}$ can be chosen arbitrarily.
Remark 1.
The time derivatives $\partial_{t}\alpha_{1}(0)$ and $\partial_{t}\alpha_{2}(0)$ also serve as initial data. In this case, we obtain the relations
$\begin{split}\nabla_{x}\bigl(\partial_{t}\Phi(0,x)\cdot\tau(0)\bigr)\big|_{x=\ell(s)}\cdot n&=\partial_{t}\alpha_{1}(0)-\kappa(0)v(0),\\ \nabla_{x}\bigl(\partial_{t}\Phi(0,x)\cdot\tau(0)\bigr)\big|_{x=\ell(s)}\cdot b&=\partial_{t}\alpha_{2}(0).\end{split}$
These formulas show that $\partial_{t}\alpha_{1}(0)$ and $\partial_{t}\alpha_{2}(0)$ are determined by the Eulerian shear of the background initial velocity field and the curvature of the particle trajectory. If the Eulerian shear vanishes at the initial time, then
$\partial_{t}\alpha_{1}(0)=\kappa(0)v(0),\qquad\partial_{t}\alpha_{2}(0)=0,$
and the subsequent transverse dynamics are driven purely by the Lagrangian geometry of the vortex curve. This shear-free setting is physically natural and theoretically significant: while strong shear can always trigger vortex instabilities, the absence of shear isolates the intrinsic geometric mechanism of vortex breakdown.
Remark 2.
We now give a definition of the occurrence of vortex breakdown at a time $t_{0}$. Define the axis of swirling particles $\zeta(t)$ by
(6) $\zeta(t):=\tau(t)-\alpha_{1}(t)\,n(t)-\alpha_{2}(t)\,b(t).$
On the other hand, the vortex axis $\zeta^{*}(t)$ is defined by
$\zeta^{*}(t):=\partial_{s}\Phi(t,s,R_{1},R_{2})\big|_{R_{1}=R_{2}=0}.$
We say that vortex breakdown occurs at time $t_{0}$ if
$\cos^{-1}\!\left(\frac{\zeta^{*}(t)}{|\zeta^{*}(t)|}\cdot\frac{\zeta(t)}{|\zeta(t)|}\right)\longrightarrow\frac{\pi}{2}\qquad(t\to t_{0}).$
Then the main theorem is as follows:
Theorem 1.
The quantities $\alpha_{1}$ and $\alpha_{2}$, which govern the occurrence of vortex breakdown, satisfy the following second–order wave equations:
$\begin{split}\alpha_{1}^{\prime\prime}(t)&=\frac{v^{\prime\prime}(t)}{v(t)}\,\alpha_{1}(t)+2v(t)\kappa^{\prime}(t)+4v^{\prime}(t)\kappa(t),\\ \alpha_{2}^{\prime\prime}(t)&=\frac{v^{\prime\prime}(t)}{v(t)}\,\alpha_{2}(t).\end{split}$
Remark 3.
In a separate line of research, Hasimoto’s remarkable discovery that the Local Induction Approximation (LIA) reduces to the focusing nonlinear Schrödinger equation revealed a profound connection between filament geometry and dispersive wave dynamics (see [4]). Despite these influential developments, existing theories rely fundamentally on vortex–filament approximations, such as concentrated vorticity, Biot–Savart expansions, LIA truncations, or asymptotic regimes in which the vortex core radius is assumed to be small. Consequently, the emergence of wave phenomena has traditionally been viewed as a feature of singular or nearly singular vortex models. It has therefore remained unclear whether similar wave dynamics arise intrinsically within the full incompressible Euler equations in a smooth, non–filament setting.
To the best of our knowledge, this paper is the first derivation showing that wave equations governing Lagrangian motion emerge directly from the incompressible Euler equations in a fully smooth setting.
2. Proof of Theorem 1
The key point of the proof is a direct computation of $\partial_{R_{1}}(\partial_{t}^{2}\Phi\cdot\tau)$ and $\partial_{R_{2}}(\partial_{t}^{2}\Phi\cdot\tau)$ along the transported curve $\bigcup_{t}\Phi(t,\ell(s))$. Let us recall the Frenet–Serret formulas:
$\frac{d}{dz}\begin{pmatrix}\tau\\ n\\ b\end{pmatrix}=\begin{pmatrix}0&\kappa&0\\ -\kappa&0&T\\ 0&-T&0\end{pmatrix}\begin{pmatrix}\tau\\ n\\ b\end{pmatrix}$.
By combining the Frenet–Serret formulas with (4) and (5), we obtain the expansions
(7) $\begin{cases}\partial_{z}\Phi=\tau+R_{1}(Tb-\kappa\tau)-R_{2}Tn\\ \qquad\quad+(R_{1}\alpha_{1}+R_{2}\alpha_{2})\kappa n+(R_{1}\partial_{z}\alpha_{1}+R_{2}\partial_{z}\alpha_{2})\tau+\mathcal{O}(R_{1}^{2}+R_{2}^{2}),\\[5.69054pt] \partial_{R_{1}}\Phi=n+\alpha_{1}\tau+\mathcal{O}(|R_{1}|+|R_{2}|),\\[2.84526pt] \partial_{R_{2}}\Phi=b+\alpha_{2}\tau+\mathcal{O}(|R_{1}|+|R_{2}|),\end{cases}$
where $\mathcal{O}(\cdot)$ denotes Landau’s notation. Equivalently, (7) can be written in matrix form as
[…]
where
$\begin{cases}A=(1-\kappa R_{1})+(R_{1}\partial_{z}\alpha_{1}+R_{2}\partial_{z}\alpha_{2}),\\ B=-R_{2}T+(R_{1}\alpha_{1}+R_{2}\alpha_{2})\kappa,\\ C=R_{1}T.\end{cases}$
We omit higher–order terms since we will eventually take the limit $R_{1},R_{2}\to 0$. The corresponding inverse matrix is
(8) $\begin{pmatrix}\tau\\ n\\ b\end{pmatrix}=\frac{1}{D}\begin{pmatrix}1&-B&-C\\ -\alpha_{1}&A-C\alpha_{2}&C\alpha_{1}\\ -\alpha_{2}&B\alpha_{2}&A-B\alpha_{1}\end{pmatrix}\begin{pmatrix}\partial_{z}\Phi\\ \partial_{R_{1}}\Phi\\ \partial_{R_{2}}\Phi\end{pmatrix}$,
where $D=A-B\alpha_{1}-C\alpha_{2}$. Moreover, differentiating (7) once more in $z$ and using the Frenet–Serret relations, we obtain
$\begin{split}\partial_{z}^{2}\Phi=&\ \kappa n-R_{1}(T^{2}+\kappa^{2})n+R_{1}\bigl((\partial_{z}T)b-(\partial_{z}\kappa)\tau\bigr)\\ &\ -R_{2}T(-\kappa\tau+Tb)-R_{2}(\partial_{z}T)n\\ &\ +(R_{1}\partial_{z}^{2}\alpha_{1}+R_{2}\partial_{z}^{2}\alpha_{2})\tau+2(R_{1}\partial_{z}\alpha_{1}+R_{2}\partial_{z}\alpha_{2})\kappa n\\ &\ +(R_{1}\alpha_{1}+R_{2}\alpha_{2})(\partial_{z}\kappa)n+(R_{1}\alpha_{1}+R_{2}\alpha_{2})\kappa(-\kappa\tau+Tb)\\ &\ +\mathcal{O}(R_{1}^{2}+R_{2}^{2}).\end{split}$
Consequently,
$\begin{split}\partial_{R_{1}}\partial_{z}\Phi\big|_{R_{1}=R_{2}=0}&=Tb-\kappa\tau+\alpha_{1}\kappa n+\partial_{z}\alpha_{1}\,\tau,\\ \partial_{R_{1}}\partial_{z}^{2}\Phi\big|_{R_{1}=R_{2}=0}&=-(T^{2}+\kappa^{2})n+\bigl((\partial_{z}T)b-(\partial_{z}\kappa)\tau\bigr)\\ &\qquad+\partial_{z}^{2}\alpha_{1}\,\tau+2(\partial_{z}\alpha_{1})\kappa n+\alpha_{1}(\partial_{z}\kappa)n-\alpha_{1}\kappa(\kappa\tau-Tb),\\ \partial_{R_{2}}\partial_{z}\Phi\big|_{R_{1}=R_{2}=0}&=-Tn+\alpha_{2}\kappa n+\partial_{z}\alpha_{2}\,\tau,\\ \partial_{R_{2}}\partial_{z}^{2}\Phi\big|_{R_{1}=R_{2}=0}&=-T(-\kappa\tau+Tb)-(\partial_{z}T)n\\ &\qquad+\partial_{z}^{2}\alpha_{2}\,\tau+2(\partial_{z}\alpha_{2})\kappa n+\alpha_{2}(\partial_{z}\kappa)n-\alpha_{2}\kappa(\kappa\tau-Tb).\end{split}$
Thus, using the identities obtained above and the orthonormality of the Frenet frame, we have
$\bigl(\partial_{R_{1}}\partial_{z}^{2}\Phi\bigr)\cdot\tau\Big|_{R_{1}=R_{2}=0}=-\partial_{z}\kappa+\partial_{z}^{2}\alpha_{1}-\kappa^{2}\alpha_{1},\qquad\bigl(\partial_{R_{1}}\partial_{z}\Phi\bigr)\cdot\tau\Big|_{R_{1}=R_{2}=0}=-\kappa+\partial_{z}\alpha_{1},$
and
$\bigl(\partial_{R_{2}}\partial_{z}^{2}\Phi\bigr)\cdot\tau\Big|_{R_{1}=R_{2}=0}=T\kappa+\partial_{z}^{2}\alpha_{2}-\kappa^{2}\alpha_{2},\qquad\bigl(\partial_{R_{2}}\partial_{z}\Phi\bigr)\cdot\tau\Big|_{R_{1}=R_{2}=0}=\partial_{z}\alpha_{2}.$
On the other hand, since $\partial_{t}=v\,\partial_{z}$ along the vortex axis, the Leibniz rule gives
$\partial_{t}^{2}\Phi=v^{2}\partial_{z}^{2}\Phi+v^{\prime}\partial_{z}\Phi,$
where $v^{\prime}=\partial_{t}v$ (and we omit writing the variables for brevity). Recall that the prime notation denotes differentiation with respect to time $t$, and that $\partial_{z}$ denotes differentiation with respect to the arc–length parameter $z$. Therefore, for $R_{1}=R_{2}=0$, we obtain
(9) $\begin{split}-\partial_{R_{1}}(\nabla p\cdot\tau)&=\partial_{R_{1}}(\partial_{t}^{2}\Phi\cdot\tau)=(\partial_{R_{1}}\partial_{t}^{2}\Phi)\cdot\tau\\ &=(-\kappa+\partial_{z}\alpha_{1})v^{\prime}+(-\partial_{z}\kappa+\partial_{z}^{2}\alpha_{1}-\kappa^{2}\alpha_{1})v^{2}\\ &=-v^{\prime}\kappa-v^{2}\partial_{z}\kappa+\partial_{t}^{2}\alpha_{1}-v^{2}\kappa^{2}\alpha_{1},\\[5.69054pt] -\partial_{R_{2}}(\nabla p\cdot\tau)&=\partial_{R_{2}}(\partial_{t}^{2}\Phi\cdot\tau)=(\partial_{R_{2}}\partial_{t}^{2}\Phi)\cdot\tau\\ &=(\partial_{z}\alpha_{2})v^{\prime}+(T\kappa+\partial_{z}^{2}\alpha_{2}-\kappa^{2}\alpha_{2})v^{2}\\ &=v^{2}T\kappa+\partial_{t}^{2}\alpha_{2}-v^{2}\kappa^{2}\alpha_{2}.\end{split}$
Next, we derive complementary identities using the Euler equations. Along the vortex axis we have $\partial_{t}\Phi=v\tau$, and hence
$\kappa n=\partial_{z}^{2}\Phi(t(z),\ell(s))=\partial_{z}\bigl(\partial_{t}\Phi\,v^{-1}\bigr)=\partial_{t}^{2}\Phi\,v^{-2}-\partial_{t}\Phi\,v^{\prime}\,v^{-3}.$
Consequently,
$\partial_{t}^{2}\Phi=v^{2}\kappa n+v^{\prime}\tau.$
Using the Euler equations $\partial_{t}^{2}\Phi=-(\nabla p)\circ\Phi$, we obtain, for $R_{1}=R_{2}=0$ (again omitting $\circ\Phi$),
$-\nabla p\cdot\tau$ $=$ $\partial_{t}^{2}\Phi\cdot\tau=v^{\prime},$
$-\nabla p\cdot n$ $=$ $\partial_{t}^{2}\Phi\cdot n=v^{2}\kappa,$
$-\partial_{z}(\nabla p\cdot n)$ $=$ $\partial_{z}(v^{2}\kappa)=v^{2}\partial_{z}\kappa+2v^{\prime}\kappa,$
$-\nabla p\cdot b$ $=$ $\partial_{t}^{2}\Phi\cdot b=0.$
By the explicit inverse formula (8) and the expansion
$\begin{split}D^{-1}=1&-\Bigl((-\kappa+\partial_{z}\alpha_{1})-\alpha_{1}^{2}\kappa-T\alpha_{2}\Bigr)R_{1}\\ &-\Bigl(\partial_{z}\alpha_{2}-(-T+\alpha_{2}\kappa)\alpha_{1}\Bigr)R_{2}+\mathcal{O}(R_{1}^{2}+R_{2}^{2}),\end{split}$
we can write
$\begin{split}\nabla p\cdot\tau=&\ D^{-1}(\nabla p\cdot\partial_{z}\Phi)-BD^{-1}(\nabla p\cdot\partial_{R_{1}}\Phi)-CD^{-1}(\nabla p\cdot\partial_{R_{2}}\Phi)\\ (\text{pull back})=&\ D^{-1}\partial_{z}(p\circ\Phi)-BD^{-1}\partial_{R_{1}}(p\circ\Phi)-CD^{-1}\partial_{R_{2}}(p\circ\Phi).\end{split}$
Hence (omitting $\circ\Phi$) we obtain
$\begin{split}-\partial_{R_{1}}(\nabla p\cdot\tau)\Big|_{R_{1}=R_{2}=0}=&\ \Bigl((-\kappa+\partial_{z}\alpha_{1}-\alpha_{1}^{2}\kappa-T\alpha_{2})\partial_{z}p-\partial_{R_{1}}\partial_{z}p\\ &\quad+\alpha_{1}\kappa\,\partial_{R_{1}}p+T\,\partial_{R_{2}}p\Bigr)\Big|_{R_{1}=R_{2}=0}\\ (\text{commuting }\partial_{R_{1}}\text{ and }\partial_{z})=&\ (-\kappa+\partial_{z}\alpha_{1}-\alpha_{1}^{2}\kappa-T\alpha_{2})\partial_{z}p-\partial_{z}\partial_{R_{1}}p\\ &\quad+\alpha_{1}\kappa\,\partial_{R_{1}}p+T\,\partial_{R_{2}}p\\ =:&\ (RHS).\end{split}$
Remark 4.
We can rephrase the commutativity of $\partial_{R_{1}}$ and $\partial_{z}$ as
$\partial_{z},\partial_{R_{1}}]:=\partial_{z}\partial_{R_{1}}-\partial_{R_{1}}\partial_{z}=0,$
where $\cdot$$\cdot$] denotes the Lie bracket. For previous studies exploiting this property, see Chan–Czubak–Y [1, Section 2.5], Lichtenfelz–Y [8], and Shimizu–Y [12]; for an earlier appearance, see Ma–Wang [9, (3.7)].
Since (along $R_{1}=R_{2}=0$, again omitting $\circ\Phi$)
$\begin{split}-\partial_{R_{1}}p&=-(\nabla p\cdot n)-\alpha_{1}(\nabla p\cdot\tau)=v^{2}\kappa+v^{\prime}\alpha_{1},\\ -\partial_{z}\partial_{R_{1}}p&=\partial_{z}(v^{2}\kappa+v^{\prime}\alpha_{1})=v^{2}\partial_{z}\kappa+2v^{\prime}\kappa+v^{\prime}\partial_{z}\alpha_{1}+\frac{v^{\prime\prime}}{v}\alpha_{1},\\ -\partial_{R_{2}}p&=-(\nabla p\cdot b)-\alpha_{2}(\nabla p\cdot\tau)=v^{\prime}\alpha_{2},\\ -\partial_{z}\partial_{R_{2}}p&=\partial_{z}(v^{\prime}\alpha_{2})=\frac{v^{\prime\prime}}{v}\alpha_{2}+v^{\prime}\partial_{z}\alpha_{2},\\ -\partial_{z}p&=-(\nabla p\cdot\tau)=v^{\prime},\end{split}$
we obtain
$\begin{split}(RHS)&=-(-\kappa+\partial_{z}\alpha_{1}-\alpha_{1}^{2}\kappa-T\alpha_{2})v^{\prime}+v^{2}\partial_{z}\kappa+2v^{\prime}\kappa+v^{\prime}\partial_{z}\alpha_{1}+\frac{v^{\prime\prime}}{v}\alpha_{1}\\ &\quad-\alpha_{1}\kappa(v^{2}\kappa+v^{\prime}\alpha_{1})-Tv^{\prime}\alpha_{2}\\ &=v^{2}\partial_{z}\kappa+3v^{\prime}\kappa+\frac{v^{\prime\prime}}{v}\alpha_{1}-v^{2}\kappa^{2}\alpha_{1}.\end{split}$
On the other hand, similarly,
$\begin{split}-\partial_{R_{2}}(\nabla p\cdot\tau)&=(\partial_{z}\alpha_{2}+T\alpha_{1}-\alpha_{1}\alpha_{2}\kappa)\partial_{z}p-\partial_{R_{2}}\partial_{z}p-(T-\alpha_{2}\kappa)\partial_{R_{1}}p\\ (\text{commuting }\partial_{R_{2}}\text{ and }\partial_{z})&=(\partial_{z}\alpha_{2}+T\alpha_{1}-\alpha_{1}\alpha_{2}\kappa)\partial_{z}p-\partial_{z}\partial_{R_{2}}p-(T-\alpha_{2}\kappa)\partial_{R_{1}}p\\ &=-(\partial_{z}\alpha_{2}+T\alpha_{1}-\alpha_{1}\alpha_{2}\kappa)v^{\prime}+\frac{v^{\prime\prime}}{v}\alpha_{2}+v^{\prime}\partial_{z}\alpha_{2}\\ &\qquad+(T-\alpha_{2}\kappa)(v^{2}\kappa+v^{\prime}\alpha_{1})\\ &=\frac{v^{\prime\prime}}{v}\alpha_{2}+v^{2}T\kappa-v^{2}\kappa^{2}\alpha_{2},\end{split}$
for $R_{1}=R_{2}=0$. Combining these identities with (9), we arrive at the coupled closure equations
(10) $\begin{split}-v^{\prime}\kappa-v^{2}\partial_{z}\kappa+\partial_{t}^{2}\alpha_{1}-v^{2}\kappa^{2}\alpha_{1}&=v^{2}\partial_{z}\kappa+3v^{\prime}\kappa+\frac{v^{\prime\prime}}{v}\alpha_{1}-v^{2}\kappa^{2}\alpha_{1},\\ v^{2}T\kappa+\partial_{t}^{2}\alpha_{2}-v^{2}\kappa^{2}\alpha_{2}&=\frac{v^{\prime\prime}}{v}\alpha_{2}+v^{2}T\kappa-v^{2}\kappa^{2}\alpha_{2}.\end{split}$
Note that $\partial_{z}\kappa=v^{-1}\kappa^{\prime}$ since $\partial_{t}=v\partial_{z}$ along the vortex axis. Rearranging (10) yields
$\begin{split}\alpha_{1}^{\prime\prime}&=\frac{v^{\prime\prime}}{v}\alpha_{1}+2v\kappa^{\prime}+4v^{\prime}\kappa,\\ \alpha_{2}^{\prime\prime}&=\frac{v^{\prime\prime}}{v}\alpha_{2}.\end{split}$
These are the desired wave equations.
Acknowledgments. The author thanks to Professors Susumu Goto, Chun Liu and Yoshikazu Giga for valuable comments. Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 24H00186.